For a first order reaction, (A) ? products, the concentration of A changes from 0.1 M to 0.025 M in 40 minutes. The rate of reaction when the concentration of A is 0.01 M is ?

Grade:12

1 Answers

Avinash
8 years ago
ln[A]/[Aº] = -kt
ln(0.025/0.1) / 40min = -k
k = 0.0346/min

ln(0.01/0.1) = -0.0346t = 66.54min
rate = k[A]
rate = 0.0346/min x 0.01M = 3.46x10^M/min